OpenStudy (anonymous):

A spherical object has a diameter of 18 cm. It has a spherical inner core with a diameter of 9 cm. What is the volume of the other layer? Use 3.14 to approximate pi and express your answer in hundredths. cm3

OpenStudy (anonymous):

\[V_{T}=V_{O}-V_{I}\]\[V_{T}=\frac{4 \pi}{3}R^{3}-\frac{4 \pi}{3}r^{3}\]\[V_{T}=\frac{4 \pi}{3}(R^{3}-r^{3})\]

OpenStudy (anonymous):

I got 2670.57

alones (alones):

I believe you're right ..i got 2670.57 too ;)